OpenStudy (anonymous):

find the exact value of csc 5pi/3

OpenStudy (anonymous):

how do i find the exact value, not degrees

OpenStudy (anonymous):

isnt it also \[-2\sqrt{3}/3\]

OpenStudy (cwrw238):

don't you want the value of the cosecant of 5pi/3 ?

OpenStudy (anonymous):

Right, didn't see that Cosecant function..my bad.

OpenStudy (cwrw238):

csc = 1 / sine

OpenStudy (anonymous):

yeah i understand how you got the 300

OpenStudy (anonymous):

do you know why its negative instead of positive?

OpenStudy (cwrw238):

sin 300 = - sin 60 = sqrt3 / 2

OpenStudy (cwrw238):



OpenStudy (cwrw238):

the sine is negative in the 4th quadrant

OpenStudy (anonymous):

oh, i thought it was suppose to be negative because csc is negative on the 4th quadrant

OpenStudy (anonymous):

actually that makes sense because im looking for1/sin

OpenStudy (cwrw238):

yea csc is negative as well because csc = 1 /sin

OpenStudy (anonymous):

then whats positive on the 4th quadrant?

OpenStudy (anonymous):

cos and sec?

OpenStudy (cwrw238):

cosine and the secant ( 1/cos)

OpenStudy (anonymous):

okay thanks

OpenStudy (cwrw238):

yw